OpenStudy (awkseacodon):

A test subject is randomly selected for a diabetes test. What is the probability of getting a subject who is not diabetic, given that the test result is negative? Find the probability using the data table.

OpenStudy (awkseacodon):

0.10 0.12 0.50 0.90

OpenStudy (awkseacodon):

@KamiBug

OpenStudy (anonymous):

i would say its 0.50

OpenStudy (kamibug):

Post the data table? :)

OpenStudy (awkseacodon):

Positive Test Negative Test Subject is diabetic 35 3 Subject is not diabetic 5 28

OpenStudy (awkseacodon):

it wasnt .50 :/ but thanks anyway @KeiraG

OpenStudy (kamibug):

We're given that the result was negative, so in the demoninator will be the total of negative tests (3+28) and the numerator will have only the negative tests from people who were not diabetic.

OpenStudy (kamibug):

So...\[\frac{ 28 }{ 3+28 } = ?\]

OpenStudy (awkseacodon):

i got 37.3

OpenStudy (kamibug):

How'd you get that? :o Divide 28 by 31 CX

OpenStudy (awkseacodon):

i put it in as 28/3+28 why would 28/31 make the answer any different? i swear math is ridiculous. anyway i failed tht question im on another one

OpenStudy (awkseacodon):

well nvm i passed the test. i only missed this question.

OpenStudy (kamibug):

Oh xP If you put 28/3+28 in a calculator, it will understand it as (28/3) + 28 So simplify the denominator first :P 28+3 = 31 28 / 31 = 0.9 :)

OpenStudy (kamibug):

YAY GOOD JOB!!! <3 *high five*

OpenStudy (awkseacodon):

math is retarded ._. ik its a need but wtf i didnt even put () F*** PEMDAS t-.-t

OpenStudy (kamibug):

XDDD

OpenStudy (awkseacodon):

anyway thanks for trying kami c: *high fives back*

OpenStudy (kamibug):

Sure cx